### v1

# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

Abstract
Information retrieval in hour-long videos presents a significant challenge, even for state-of-the-art Vision-Language Models (VLMs), particularly when the desired information is localized within a small subset of frames. Long video data presents challenges for VLMs due to context window limitations and the difficulty of pinpointing frames containing the answer. Our novel video agent, FALCONEye [^1] Webpage: https://cplou99.github.io/FALCONEye/, combines a VLM and a Large Language Model (LLM) to search relevant information along the video, and locate the frames with the answer. FALCONEye novelty relies on 1) the proposed meta-architecture, which is better suited to tackle hour-long videos compared to short video approaches in the state-of-the-art; 2) a new efficient exploration algorithm to locate the information using short clips, captions and answer confidence; and 3) our state-of-the-art VLMs calibration analysis for the answer confidence. Our agent is built over a small-size VLM and a medium-size LLM being accessible to run on standard computational resources. We also release FALCON-Bench, a benchmark to evaluate long (average $>1$ hour) Video Answer Search challenges, highlighting the need for open-ended question evaluation. Our experiments show FALCONEye’s superior performance than the state-of-the-art in FALCON-Bench, and similar or better performance in related benchmarks.
Figure 1: Overview of our meta-architecture FALCONEye designed to Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs. A VLM pre-process the video generating captions from small video clips. A LLM iteratively refines the search by focusing on the most promising clips using the captions. The exploration follows a hierarchical reasoning: initially, a broad set of low-resolution frames is analyzed by the VLM for each promising clip, progressively narrowing down to fewer frames with higher resolution as the search concentrates on smaller temporal clips. The LLM uses captions, question semantics, answer completion, and confidence scores to determine whether to continue exploring or end the exploration. Once a high-confidence answer is found, or the search reaches a predefined threshold, FALCONEye outputs the final answer, its confidence score, and the corresponding temporal interval.
1 Introduction
Vision-Language Models (VLMs) have shown remarkable performance in answering complex, fine-grained questions about images and short videos [17, 4, 32], but solving similar tasks on long videos remains an open problem [5, 53]. Besides the intrinsic problem complexity, processing and encoding all video frames detailed information is too costly and impractical [54]. Our work emphasis is on analyzing very long videos that contain the answer on a tiny set of frames, a needle in a haystack problem, that we will designate as Video Answer Search (VAS) task. Thus, the challenge is to locate the specific content necessary to answer the question in an efficient and practical way.
Large-Language Models (LLMs), and consequently VLMs, are constrained by their context window size, which limits the number of tokens they can process in a single pass. This constraint is not critical to process a 300-page book or an image. However, when dealing with videos, only a small subset of frames (typically in the order of dozens or hundreds) can be sampled and encoded to keep the total number of tokens within the context window size [52, 47]. This sampling strategy may work for short videos, but is likely to fail for extremely long videos, where using only a few frames is likely to miss a significant part of relevant information. Therefore, research is increasingly focused on adapting VLMs to work with long-form videos [37, 38, 6].
First, Video Question Answering (VQA) benchmarks are extending video durations and question complexity [5, 3, 42]. However, the focus is mainly on multiple-choice questions, where the LLM’s ability to eliminate options may be even more crucial than its ability to infer the answer from the sampled frames  [35]. In this work, we release FALCON-Bench, specifically designed to evaluate VAS in (90 publicly available) long-form videos, focusing on (more than 580) open-ended questions that require to pinpoint specific information, measuring both accuracy and localization.
Second, several VLMs specifically designed for long videos have emerged [38, 37, 4], leveraging smart frame sampling and token compression techniques to maximize the amount of sampled visual information that fits within its context window. However, they often lack the precision required for VAS task when only a tiny fraction of the video contains the answer, as they provide a global representation instead of bootstrapping resources in the most relevant video segments. In this context, we introduce FALCONEye, a video agent to tackle VAS task. Through an iterative approach that integrates a VLM and a LLM, FALCONEye leverages captions, question-answers semantics, and confidence scores to optimize localization, avoiding irrelevant clips and concentrating resources on exploring temporally relevant segments until finding the answer (Figure 1).
Overall, our work presents the following contributions: 1) FALCON-Bench Benchmark: The first benchmark specifically designed for the VAS task. Unlike other VQA benchmarks, it focuses on open-ended questions that require pinpointing answers within tiny temporal regions of one-hour-long videos. 2) Meta-architecture for VAS (FALCONEye): A system that integrates VLM and LLM capabilities to emulate human-like VAS behavior. It leverages a novel search algorithm that reasons to iteratively focus on video clips most likely to contain the answer. 3) Evaluation of state-of-the-art VLMs calibration: A study assessing the calibration of leading VLMs in VQA tasks. Both benchmark and model have been built under the lmms-eval framework [51] in order to ease the reproducibility of the results and future evaluations with other models and benchmarks. FALCONEye and FALCON-Bench represent a significant advancement in VAS and, VQA in general, setting a new standard for precision-guided exploration in long-form video answer search task.
2 Related work
Table 1: Comparison of our FALCON-Bench and other popular benchmarks for VQA task in videos. We focus on labeling, number (#) of videos and question (QAs), its average and total duration in seconds (s) and hours (h), respectively, and the theme of the videos. Lastly, we analyze if the evaluation is oriented to multiple-choice (MC) or open-question (O), if the benchmark contains questions focused on single detail (SD) aspects of the video, and, if the ground truth (GT) answer contains the GT time interval in which the answer is contained.
Benchmark Labels #Videos #QAs Avg. Dur. (s) Total Dur. (h) Theme Category Eval. SD GT TIME
ActivityNet-QA [49] Human 5,800 58,000 180 40 Everyday Life O
NeXT-QA [46] Human 5,440 52,044 44 12.22 Everyday Life MC/O
MVBench [21] Auto 4,000 4,000 16 17.78 Life, Human Action, Movie MC
EgoSchema [28] Auto 5,031 5,031 180 251.55 Life, Human Action, Egocentric MC
MovieChat-1K [38] Human 130 1,950 500 18.06 Movie MC/O
HourVideo [5] Human 500 12,976 2,742 380 Life, Human Action, Egocentric MC
Event-Bench [14] Auto+Human 1294 2,190 101 36.46 Life, Egocentric, Fun, Cartoon MC (21%)
LONGVIDEOBENCH [44] Human 3,763 6,678 473 494.49 Life, Movie, Knowledge, News MC (30%)
MLVU [53] Human 1,334 2,593 720 267.8 Movie, Surveillance, Egocentric MC/O (64%)
Video-MME [15] Human 900 2,700 1,017 254.47 TV, Sports, Artistic, Life MC (15%)
LVBench [42] Human 103 1,600 4,101 117.3 Sports, TV shows, animations MC (14%)
InfiniBench [3] Auto 1219 108,200 3,155 1,068.3 Movie, TV shows MC/O (14%)
FALCON-Bench (Ours) Human 92 576 4,738 121.1 Soccer, Movie, Egocentric MC/O (100%)
Vision Language Models (VLMs).
VLMs emerged to extend LLMs’ capabilities to vision, with the seminal work of LLaVA [25], which integrated a pre-trained CLIP visual encoder, ViT-L/14 [33], to extract visual tokens, with an LLM (Vicuna [8]) for multimodal text and image understanding. This integration appears in other open-source architectures which improved the alignment and processing of visual features, such as BLIP-2 [20], LLaVA-1.5 [24], mPLUG-Owl2 [47], and LLaMa3.2-Multimodal [1]. More recently, Qwen2-VL [41] introduced a VLM with a ViT capable of processing images at arbitrary resolutions, dynamically converting them into a variable number of visual tokens. Notably, many companies have integrated increasingly powerful VLMs into their proprietary long-context chatbot systems, including OpenAI’s GPT-4V [32], Google DeepMind’s Gemini [11] or Anthropic’s Claude [2]. VLMs excel at image questions, but extending these capabilities to videos remains an open challenge due to the large number of tokens required to represent video content. VLM strategies for video extends VLMs to process multiple frames as input [22, 19, 21, 17, 52] employing a uniform video sampling strategy to extract and process up to 64 video frames, to keep the amount of visual tokens within the LLM’s context window. More specific long-video solutions target more efficient and sophisticated video representations. MovieChat [38] combines short- and long-term token-based memories to build a global video representation, LongVU [37] presents a spatiotemporal adaptive compression to retain the most relevant parts for question answering, and LongVILA [6] introduces Multi-Modal Sequence Parallelism to extend the number of video frames processed from 8 to 2048. Apollo [54] conducts an extensive study on strategies to enhance long-video understanding. Recently, Qwen2.5-VL [4] extends its dynamic resolution capabilities to the temporal dimension, allowing processing up to 768 frames. Unlike these methods focused on achieving a global understanding of the entire video, our approach exploration-based strategy bootstraps resources in the most relevant video segments for the question.
VQA Benchmarks.
Existing Video Question Answering (VQA) benchmarks evaluate VLMs capabilities for a wide range of video comprehension tasks. Early benchmarks focus on short videos and questions to understand temporal actions [46], human actions [49], egocentric videos [28] or movies [21]. Similar to ours, Event-Bench [14] focuses on single-detail events occurred within short temporal windows but uses one-minute-long videos. Recently, several VQA benchmarks [5, 44, 53, 15] have extended the temporal duration of videos to around 10 minutes. Unlike these benchmarks, which primarily focus on Multiple-Choice Questions (MCQs), MovieChat-1K [38] is centered on open-ended questions (OQs). LVBench [42] and InfiniBench [3] further push this challenge with, up to one hour, longer videos. Some of these later benchmarks include a percentage of single-detailed questions, where the answer is found within a short time window of the video, but only LONGVIDEOBENCH [44] and InfiniBench evaluate a small fraction of these questions in open-ended mode. Our FALCON-Bench is designed to address the challenges of single-detail questions and open-ended responses in very long videos. This is the first benchmark to provide ground truth time intervals that contain the answers, ensuring that models can be evaluated to find and localize answers. Table 1 presents a comparison of all these benchmarks.
Meta-architectures over VLMs for VQA.
Many strategies have emerged to tackle the most challenging VQA scenarios, often defined as meta-architectures agnostic to specific models [45]. Long Story Short [10] leverages a LLM that summarizes the video through clip captions, selects the most relevant clips, generates potential answers from their captions, and combines the likelihood of the option letter tokens with the similarity between question and frame embeddings to determine the final response. VideoAgent [43], which shares some similarities with our approach, introduces a three-step algorithm within a training-free meta-architecture. It integrates a VLM to generate clip captions and an LLM to produce answers with confidence scores, selecting the most relevant frames based on the similarity between CLIP embeddings of the frames and the question features. MoReVQA [29] and ProViQ [9] proposes a decomposed multi-stage, modular reasoning framework based on ViperGPT [39], a library of API modules to assemble executable programs for video reasoning. Ranasinghe et al. [34] present the MVU framework, which injects object-centric information—extracted via object detection and tracking models—into an LLM. TimeChat [36] fine-tunes a VLM built on LLaMa-2 by introducing two novel components: a timestamp-aware frame encoder and a sliding video Q-Former, enhancing fine-grained temporal understanding. SeViLA [48] follows a sequential approach, chaining a pre-trained image-language model (BLIP-2) to first localize relevant frames and then answer questions based on them. As most approaches are designed for short videos, longer VQA benchmarks [5] often adopt the Socratic method [50], where an LLM answers questions using captions pre-extracted by a VLM from short clips.
3 FALCONEye
We introduce FALCONEye, a meta-architecture that combines a VLM and an LLM within an exploration-based search algorithm to tackle the VAS task. Unlike previous agents designed for MCQs and short videos, operating at the frame level, FALCONEye is built to handle OQs and one-hour-long videos, operating at the clip level.
Figure 2: FALCONEye exploration algorithm that follows our FALCONEye meta-architecture. Given a question ( Refer to caption) and a video, it starts with a global Pre-processing of the video, where the VLM generates short captions ( Refer to caption) from video clips. Afterwards, the LLM Reasons from the Q and the captions to select some candidate clips to explore. The VLM Evaluates the candidate clips sampling frames varying number and resolution according to the clip duration. We get an answer ( Refer to caption) and its confidence ( Refer to caption) for each candidate clip. From all { Refer to caption, Refer to caption, Refer to caption} tuples, the LLM Decides wether the final answer has already been found or, otherwise, it returns some promising clips to continue exploring. If these promising clips are large enough, we segment them to generate new shorter candidate clips to evaluate. If the promising clips are already too short ( Refer to caption), we ask again the LLM to reason removing the caption ( Refer to caption) of the already explored clip.
Problem definition:
Given a video-question pair ($\mathcal{V}$, $Q$). The goal of the VAS task is:
  1. 1.
    Find the answer $\mathcal{A}$ that accurately responds to $Q$.
  2. 2.
    Localize a short video clip $v\in\mathcal{V}$ as evidence, ensuring that the answer $\mathcal{A}$ can be observed on its frames.
3.1 Meta-architecture
We propose a meta-architecture with two main components: a Vision-Language Model (VLM) and a Large Language Model (LLM). This meta-architecture is agnostic to any specific LLM or VLM, allowing flexibility in model selection, including frozen foundation models. Our overall idea involves using a lightweight VLM to process the most computationally expensive data (visual information) and a medium-sized LLM to handle text-based reasoning, which is significantly cheaper to process. This setup enables higher efficiency and cost reduction compared to addressing the task with current large-scale state-of-the-art VLMs. The VLM is used only to process short video clips, namely for two tasks, provide captions and generate potential answers to $Q$. The LLM powerful reasoning capabilities are exploited to extract information from all the captions and find interesting clips, and decide wether to explore other clips, refine the current clip search (zooming into it) or if the final answer has been found.
3.2 Exploration Algorithm
The exploration algorithm consists of four main stages: 1 Pre-processing, 2 Reasoning, 3 Evaluation, and 4 Decision (Figure 2). First, a pre-processing phase, which can be done offline for any video as it is question independent. For question-answering, stages 2- 4 are executed in a loop, which ends when either the LLM selects a potential answer as the final one, the LLM reasons that there are no more candidate clips to explore, or a maximum number of iterations is reached. Algorithm 1 and Section 9.2 in the supplementary material detail the pseudo-code of this algorithm and all the prompts used to interact with the LLM.
1 Pre-processing-. Here, our method generates short captions and a concise summary of the video. The video $\mathcal{V}$ is segmented into short video clips $\{v_{1},\ldots,v_{n}\}$, that we call first-level clips, such that $\mathcal{V}=\cup_{i=1}^{n}v_{i}$. For each video clip $v_{i}$, the VLM uniformly samples a small set of frames to generate a short caption $c_{i}$ that summarizes its content. Given the set of all captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$, the LLM then generates a global yet brief summary $\mathcal{S}$ of the video.
2
Reasoning-. Given the summary $\mathcal{S}$, the set of captions $\mathcal{C}$, and the question $Q$, the LLM pinpoints a set of first-level clips $\{v^{*}_{1},\ldots,v^{*}_{m}\}\subset\mathcal{V}$ as candidates to contain the answer. The number of selected clips $0\leq m\leq n$ is directly selected by the LLM and it varies each iteration. If the LLM determines that no clip is likely to contain the answer, i.e.: $m=0$, the process terminates.
3
Evaluation-. In this stage, each candidate clip $v^{*}_{i}$ is evaluated by the VLM. First, a uniform sampling of frames is performed within the clip. The number and resolution of sampled frames vary according to the duration of $v^{*}_{i}$, creating a zoom-in effect. For first-level clips, a larger number of frames with lower resolution are sampled, whereas for shorter clips, fewer frames with higher resolution are selected. For fixed input resolution VLMs, we use tiling to increase the resolution (similar to what GPT4o [32] does).
The VLM takes as input the question $Q$ along with the sampled frames from $v^{*}_{i}$ and generates a potential answer $\mathcal{A}^{*}_{i}$ and a confidence score $p(\mathcal{A}^{*}_{i})$. The answer is derived from a set of generated tokens $\mathcal{T}\in D$, such that $\mathcal{A}^{*}_{i}=Tokenizer(\mathcal{T})$, where $D=\{t_{1},\ldots,t_{d}\}$ is the token dictionary used by the LLM. During the VLM generation process, we store the logits $\{l_{1}^{k},\ldots,l_{d}^{k}\}$ for each generated token $t_{k}\in\mathcal{T}$ and compute $p(\mathcal{A}^{*}_{i})$ as the geometric average of the answer token probabilities [26]:
$p(\mathcal{A}^{*}_{i})=\sqrt[s]{\prod_{k=1}^{s}\max\left(Softmax\left(\{l_{1}^{k},\ldots,l_{d}^{k}\}\right)\right)},$ (1)
where $s$ is the number of generated tokens in $\mathcal{T}$ (see Fig. 9 in supp. material). In MCQ scenario, there is no need for an average as $s=1$ (a single token for the option letter: A, B, C, D, …) and the confidence is just the softmax of the logits for the option letter token [10]. This step is supported by our VLM calibration analysis presented later in Section 6.
4
Decision-. At this stage, the LLM receives the information of the candidate clips that have just been evaluated in the last iteration of the Evaluation stage, $\{\mathcal{A}^{*}_{i},p(\mathcal{A}^{*}_{i}),c_{i}\}$, along with the summary $\mathcal{S}$ and the question $Q$. Based on confidence values, answer quality, captions, and the temporal localization of the clips, the LLM determines whether one of these answers should be selected as the final answer. If an answer meets the selection criteria, the process returns the final answer $\mathcal{A}$, its temporal localization, and its confidence $p(\mathcal{A})$. Otherwise, the system identifies which of the candidate clips $\{v^{*}_{1},\ldots,v^{*}_{m}\}$ remain promising for further exploration and zoom-in. The number of promising clips is not fixed and varies based on the LLM’s judgment. If the identified promising clips are first-order or second-order, they are segmented into shorter clips, converting them into second-order and third-order clips, respectively. These new clips form an updated set of candidate clips $\{v^{*}_{1},\ldots,v^{*}_{m}\}\subset\mathcal{V}$, and the process returns to Stage 3 Evaluation. If the selected promising clips are at the third-order level, the system returns to Stage 2 Reasoning. To prevent redundant exploration, captions corresponding to previously evaluated first-order clips are removed from the caption set $\mathcal{C}$.
If the loop between stages 2 and 4 ends without producing a final answer, a final decision must be made. This situation occurs when either the maximum number of iterations—defined as the number of candidate clips evaluated by the VLM—has been reached, or when the LLM determines that no more first-order clips remain as candidates for exploration. In this case, the LLM gathers the information from all previously evaluated clips $\{\mathcal{A}^{*}_{i},p(\mathcal{A}^{*}_{i}),c_{i}\}$ and selects the most suitable answer among them. Finally, the process outputs the final answer $\mathcal{A}$, its temporal localization, and its confidence value $p(\mathcal{A})$.
4 FALCON-Bench
The proposed benchmark is specifically designed for the VAS task, using one-hour-long videos obtained from external public datasets. Our contribution lies in the constructed set of Q&A that assess the VAS task on these videos. Each question is carefully crafted to ensure a unique answer within the video, with the answer content embedded in a single, short temporal window. Unlike previous VQA benchmarks, which often rely on broad video-level reasoning, we highlight the challenge of finding a needle in a haystack within long-form videos. Our benchmark emphasizes the necessity of not just uniformly sampling frames or applying token compression across the entire video, but instead reasoning over the content to localize the exact clip that contains the answer. To facilitate the evaluation of the localization subtask, we provide the ground-truth (GT) temporal interval where the answer is contained.
The videos of our benchmark were sourced from SoccerNet [16], MovieChat-1K [38] and Walking Tours Dataset [40], ensuring diversity and comprehensive coverage across different video types. Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Each question is accompanied by four different answer options, allowing for multiple-choice question (MCQs) evaluation. However, the primary focus remains on open-ended questions (OQs), as this approach is more aligned with the VAS task. The annotation process was conducted by humans assisted by GPT-4o for generating answer options. More details on the benchmark content and construction process are in Sec. 8 of the supplementary material. The benchmark has been built under the lmms-eval framework [51] to ease further evaluation and will be publicly available upon acceptance.
Table 2: Models performance comparison in FALCON-Bench test split, both MCQs and OQs. We provide LLM size (S), mean number of frames (#F) processed per question by each VLM. Results grouped by MovieChat (M), Soccer (S), WalkingTours (T), and average (Avg.).
Model Multiple-Choice Questions Open Questions
Name Config. Time Accuracy Loc. Time Accuracy Score Loc.
S #F s M S T Avg. mGToU s M S T Avg. M S T Avg. mGToU
Baselines
Full Mark - - - 100 100 100 100 100 - 100 100 100 100 5.00 5.00 5.00 5.00 100
Human - - 115.6 100 100 76.9 92.3 73.3 129.2 84.6 95.0 76.5 85.4 4.31 4.80 4.06 4.42 69.2
GPT-4o-mini-blind N/A - 0.48 45.0 32.4 42.0 39.8 0.00 1.76 8.82 0.00 2.00 3.60 0.48 0.00 0.18 0.22 0.00
Qwen2.5-VL-blind 7B - 0.08 42.0 34.1 48.0 41.4 0.00 2.45 1.96 0.28 2.00 1.41 0.21 0.11 0.30 0.20 0.00
Random - - 0.00 25.0 25.0 25.0 25.0 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Open-Source VLMs with Multi-Image Support
mPLUG-Owl3 [47] 7B 16 5.52 41.1 28.5 36.0 35.2 N/A 6.63 8.82 4.23 10.0 7.68 0.68 0.29 0.66 0.54 N/A
LLaVA-v1.5 [25] 7B 32 11.4 23.5 31.9 44.0 33.1 N/A 12.7 3.92 4.80 2.00 3.57 0.36 0.31 0.24 0.30 N/A
LLaVA-v1.6 [19] 7B 32 12.9 28.4 30.5 38.0 32.3 N/A 14.3 4.90 4.51 2.00 3.80 0.31 0.30 0.20 0.27 N/A
LLaVA-OV [17] 7B 32 16.9 48.0 30.2 46.0 41.4 N/A 7.29 9.80 4.23 16.0 10.0 0.53 0.25 1.00 0.59 N/A
Qwen2.5-VL [4] 7B 768 68.9 50.0 24.4 48.0 40.8 N/A 70.1 11.7 8.19 10.0 9.96 0.72 0.53 0.62 0.62 N/A
Open-Source VLMs designed for videos
Video-LLaVA [22] 7B 8 3.37 28.4 26.8 34.0 29.7 N/A 4.50 6.86 5.08 8.00 6.64 0.56 0.37 0.68 0.53 N/A
VideoChat2-HD [21] 7B 8 6.74 29.4 24.2 26.0 26.5 N/A 8.06 8.82 2.82 8.00 6.54 0.71 0.20 0.52 0.47 N/A
LLaVA-Video [52] 7B 32 12.4 50.0 38.9 58.0 48.9 N/A 13.0 10.7 11.0 12.0 11.2 0.76 0.66 0.86 0.76 N/A
Open-Source VLMs specific for long videos
MovieChat-OV [38] 7B 1024 62.3 43.1 38.7 46.0 42.6 N/A 69.3 9.80 4.80 10.0 8.20 0.53 0.25 0.66 0.48 N/A
Apollo [54] 7B 0.5fps 36.2 51.9 35.8 64.0 50.5 N/A 36.5 16.6 10.7 18.0 15.1 0.96 0.68 1.06 0.90 N/A
Meta-architectures built from a LLM (GPT4o-mini) and a VLM (Qwen2.5-VL)
Socratic [50] 7B+N/A 0.5fps 186.6 53.9 48.0 54.0 51.9 21.3 186.7 25.4 29.0 20.0 24.8 1.49 1.71 1.16 1.45 19.6
VideoAgent++ [43] 7B+N/A 9.2 45.0 48.0 52.2 36.0 45.4 19.1 43.8 8.82 24.8 8.00 13.8 0.56 1.38 0.44 0.79 13.3
FALCONEye [Uncaptioned image] 7B+N/A 659.2 393.0 74.5 61.5 74.0 70.0 27.7 245.4 50.9 37.2 46.0 44.7 2.67 1.96 2.50 2.38 24.9
FALCONEye [Uncaptioned image] 7B+N/A 518.5 213.3 67.6 58.4 68.0 64.7 25.2 187.9 43.1 38.4 42.0 41.1 2.36 2.03 2.28 2.22 22.7
N/A means Not Available, due to unknown model size or capability not available in the approach. – means Does not apply.
5 Experiments
Metrics-.
In MCQs, we report accuracy as the primary evaluation metric. For OQs, following the approach of Maaz et al. [27], also applied in others, as InfiniBench [3], we employ GPT-assisted evaluation. GPT reports accuracy (100 or 0) and a score value (0-5 range), assessing key aspects such as Correctness of Information, Detailed Orientation, Contextual Understanding, Temporal Understanding and Consistency. We also use the Ground Truth over Union (GToU) metric to evaluate the quality of the predicted temporal interval, as a visual evidence (short clip) that contains the answer. Section 8.3 of the Supplementary discusses why we find GToU more adequate than mIoU (a clip that tries to match the entire GT clip).
Table 3: Meta-architectures cost and performance comparison in the test split of the FALCON-Bench OQs. We set LLM=GPT-4o-mini, VLM=Qwen2.5-VL. We measure LLM usage as number of inferences (#Inf.) per question and, total tokens (#Toks) and time (s) per inference. We measure VLM usage as number of inferences (#Inf.T) per question, and number of frames (#F), max output tokens (#T) and time (s) per inference. We compute #Inf.T adding the #Inf per Q and per V as #$\#Inf.T=\#Inf.Q+\#Inf.V/Q.V$, knowing that $Q.V=6.24$. VLM inferences can be for Captioning or VQA. We report the average accuracy (Acc.), score (Sc.) and localization (Loc.).
LLM VLM for Captioning or VQA Total Performance
Usage Avg. per Inf #Inf. Avg. per Inf Time Open Questions
Model Name Mode #Inf. #Toks s V Q T #F #T s s Acc. Sc. Loc.
Socratic [50] QA 1 13.5K 2.1 78.9 - 12.6 30 768 14.6 186.7 24.8 1.45 19.6
Socratic [Uncaptioned image] [50] QA 1 2.4K 1.1 78.9 - 12.6 30 64 9.8 125.0 15.5 1.03 18.3
VideoAgent++ [43] QA 5.6 1.3K 3.2 5 2.90 3.70 1 64 2.2 43.8 13.8 0.79 13.3
SequentialBP - - - - - 53.3 53.3 120 64 12.7 676.9 28.7 1.52 13.7
Sequential - - - - - 78.9 78.9 120 64 12.7 1002 34.8 1.83 9.05
IterativeSamp - - - - - 85.4 85.4 120 64 18.7 1596 24.0 1.34 1.26
FALCONEye [Uncaptioned image] Reason 4.3 1.7K 5.8 78.9 9.98 22.5 29.3 64 9.8 245.4 44.7 2.50 24.9
FALCONEye [Uncaptioned image] Reason 3.0 1.7K 5.8 78.9 4.86 17.4 29.8 64 9.8 187.9 41.8 2.24 22.7
Illustration of exploration baselines
Exploration Baselines-.
A widely used baseline in recent long VQA benchmarks [5] is the Socratic method [50]. In this baseline, a VLM generates captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$ from short video clips $\{v_{1},\ldots,v_{n}\}\in\mathcal{V}$ and a LLM generates $\mathcal{A}$ from $Q$ and $\mathcal{C}$. In addition to this baseline, we introduce three alternative exploration baselines (Fig. within Tab. 3) that rely solely on a VLM, guided by its answer confidence computed as described in Sec 3.
  1. 1.
    Sequential: it sequentially processes the entire video $\mathcal{V}$ in short clips $\{v_{1},\ldots,v_{n}\}$. For each short clip $v_{i}$, the VLM generates a potential answer $\{\mathcal{A}^{*}_{i},p(\mathcal{A}^{*}_{i})\}$. At the end of the process, the final answer is selected as $A=\mathcal{A}^{*}_{k}$, where $k=\arg\max_{i}p(\mathcal{A}^{*}_{i})$.
  2. 2.
    SequentialBP: similar to the Sequential baseline, but with an early stopping criterion. The process halts as soon as an answer $\mathcal{A}^{*}_{i}$ is found with a confidence score $p(\mathcal{A}^{*}_{i})$ exceeding a predefined threshold.
  3. 3.
    IterativeSampling: it iteratively generates potential answers within a progressively refined video clip $v^{*}$. The search begins with $v^{*}=\mathcal{V}$. The algorithm selects a small set of frames around the most unexplored regions within $v^{*}$ and the VLM generates $\{\mathcal{A}^{*},p(\mathcal{A}^{*})\}$. Whenever a higher-confidence answer is found, $v^{*}$ is updated to the temporal window where the sampled frames were located, further narrowing the focus.
FALCONEye-.
We evaluate our meta-architecture using Qwen2.5-VL [4] as the VLM and GPT4o-mini [31] as the LLM. As detailed in Sec. 6, Qwen2.5-VL outperforms other small-size VLMs on FALCONBench OQs in short clips near the answer, and is well calibrated. Its ViT supports images of any resolution, dynamically adjusting visual tokens, enabling an accurate zoom-in effect as clip lengths decrease. GPT4o-mini provides a balanced trade-off between cost and reasoning power (Sec. 6). Prompts used for LLM queries are detailed in Sec. 9.2. Inspired by the Socratic approach [5], first-level clips are 1min length, but captions are now shorter formed by up to 64 tokens (instead of 768). Within these clips, second-order (20s) and third-order (5s) clips are explored. The zoom-in effect, validated in Sec. 6, adjusts from 30 frames ($672\times 1204$) in first-level clips to 10 frames ($824\times 1462$) in third-order clips. In OQs, the LLM determines the final answer based on completion, confidence ($0.8$ threshold), captions, and temporal localization (as explained in Sec. 3). In MCQs, this evaluation is replaced by a simple confidence-based check with $0.9$ threshold since all possible answers are predefined (A, B, C, or D) and inherently valid, making confidence the only selection criterion. Lastly, we evaluate two versions of FALCONEye: FALCONEye [Uncaptioned image], the top-performing configuration, which evaluates up to 45 candidate clips, and FALCONEye [Uncaptioned image], a cost-efficient variant that limits evaluation to 10 candidates.
Table 4: Model performance comparison in LongVideoBench and MLVU benchmarks.
Model MLVU tasks LongVideoBench tasks *
Name S # F NQA ER PQA S2E S2O S2A E2O O2E
858s 575s 466s 727s 716s 642s 470s 672s
Baselines
Full Mark - - 100 100 100 100 100 100 100 100
Random - - 16.7 16.7 16.7 20.0 20.0 20.0 20.0 20.0
Proprietary Long-Context LLMs
Gemini-1.5-Pro N/A 256 N/A N/A N/A 74.6 58.3 76.2 68.7 73.3
GPT-4o N/A 0.5fps 42.9 47.8 57.1 76.8 69.8 70.9 67.3 72.8
Open-Source Image LLMs with Multi-Image Support
LLaVA-OV [17] 72B 32 46.7 58.4 58.0 N/A N/A N/A N/A N/A
InternVL2.5 [7] 72B 64 48.3 47.2 52.0 N/A N/A N/A N/A N/A
LLaVA-Video [52] 34B 32 N/A N/A N/A 62.1 50.2 51.2 50.9 58.5
Open-Source Video LLMs
VILA-1.5 [23] 40B 14 38.3 35.8 62.0 N/A N/A N/A N/A N/A
Aria [18] 25B 256 56.6 66.0 58.3 N/A N/A N/A N/A N/A
Meta-architectures built from a LLM (GPT4o-mini) and a VLM (Qwen2.5-VL)
FALCONEye [Uncaptioned image] 7B+N/A 2fps 61.7 43.5 62.9 60.2 43.1 64.8 60.0 64.4
FALCONEye [Uncaptioned image] 7B+N/A 2fps 62.0 40.9 59.6 59.1 38.9 63.6 53.8 62.1
– : Does not apply. N/A: Not Available due to unknown model size or not published results on benchmark.
* Our results are in validation split since gt test split is not publicly available.
5.1 FALCON-Bench results
Table 2 presents the performance of baselines, open-source VLMs, and meta-architectures on FALCONBench. As baselines, we include human performance both to establish an upper bound and to analyze human strategies in VAS, which guided our algorithm design (Sec. 8.4). Additionally, we evaluate our two selected LLMs (GPT4o-mini and Qwen2.5-VL) in blind mode (without visual information) to assess their reliance on text-based priors. For open-source VLMs, we test small-scale versions of the main state-of-the-art models, including Apollo and MovieChat, which are specifically designed for long videos. Additionally, we evaluate two well-known meta-architectures, Socratic [50] and VideoAgent [43]. To ensure a fair comparison with our FALCONEye, both meta-architectures are implemented using the same VLM (Qwen2.5-VL) and LLM (GPT4o-mini). The Socratic approach generates captions (768 tokens) from 1-minute clips, while VideoAgent is extended to handle open-ended questions and predict answer localization, resulting in VideoAgent++. Most VLMs that are limited to a small number of sampled frames fail to outperform the LLM-blind baselines in MCQs, suggesting that their advantage over random guessing stems from discarding certain options, rather than actual answer retrieval from visual content. A similar issue arises with Qwen2.5-VL when operating in its default mode (2fps, up to 768 frames), as it downscales frames to $140\times 240$, losing crucial details. Only LLaVA-Video and Apollo show clear improvements over the blind baselines. This trend is not observed in OQs, where models must generate answers independently, validating their reliability. Among meta-architectures, only FALCONEye demonstrates robustness for long-form VAS, achieving 70.0% accuracy in MCQs and 44.7% in OQs with its top-performance configuration, and 64.7% and 41.1%, respectively, in the cost-efficient variant. The latter configuration achieves performance close to the Socratic baseline while maintaining a processing time per question not so far from human response times. We analyze their costs and performances, including exploration baselines, in Table 3. We also include Socratic [Uncaptioned image], which utilizes the short captions (64 tokens) generated by FALCONEye. While baseline exploration algorithms outperform existing meta-architectures, they require significantly more time. Notably, our LLM-guided exploration algorithm reduces this overhead while simultaneously improving performance by efficiently narrowing the search.
5.2 Other benchmarks results
We evaluate FALCONEye on additional VQA benchmarks to compare its performance against state-of-the-art private and public large-scale VLMs. From the benchmarks listed in Tab. 1, we select MLVU [53] and LONGVIDEOBENCH [44], as both were recently published in top-tier conferences, include tasks related to single-detail questions (closely aligned with VAS), and are available within the lmms-eval framework. A summary of these results is presented in Tab. 4.
Although these benchmarks do not fully capture the two key challenges addressed by FALCONEye: one-hour-long videos (these benchmarks instead average only a few minutes per video, where the 256 frames sampled by top VLMs represent a significant portion of the entire video), and open-ended question (OQ) evaluation (these benchmarks primarily assess MCQs)—FALCONEye demonstrates competitive performance against top VLMs, achieving the new state-of-the-art in some MLVU subtasks.
Table 5: Some VLMs calibration and performance comparison with GT 1min-length clips in the test split of FALCONBench.
Model Visual Information MCQs OQs
Name $\Delta t$ Frames Acc. Calib. GPT Calib.
s # Res. Avg. ACE $\downarrow$ CC $\uparrow$ Acc. Sc. ACE $\downarrow$ CC $\uparrow$
LLaVA-Video 60 32 $336\times 336$ 79.2 0.11 0.35 52.6 2.81 0.19 0.06
LLaVA-OV 60 32 $336\times 336$ 72.2 0.35 0.0 39.8 2.15 0.27 5e-3
Qwen2.5-VL 60 2fps $336\times 616$ 73.1 0.12 0.22 59.5 3.24 0.29 0.28
Figure 3: Qwen2.5-VL calibration plots with the GT 1min-length clips of the FALCONBench test split. Left: MCQs. Right: OQs.
6 Configuration Analysis
VLMs calibration study.
Since our exploration algorithm relies on answer confidence, here we evaluate if the confidence values are well calibrated. For that, we adopt Reliability Diagrams [12], which group predictions into bins based on confidence and measure the gap (calibration error) between confidence and accuracy for each bin. From these plots, we compute the Average Calibration Error (ACE) and Calibration Count (CC), quantifying the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE).
Given the poor performance of state-of-the-art VLMs on FALCONBench, we ease the task by providing directly one-minute clips taken from the middle of the ground-truth (GT) temporal interval (Sec. 10.1 of supp. material). For MCQs, confidence computation is straightforward, as only a single token is output (Sec. 3.2). However, for OQs, confidence must be aggregated across the entire sequence of tokens. Similar to calibration studies for LLMs [26], we investigate various aggregation metrics in Sec. 10.2, identifying the geometric average (Eq.1) as the most suitable approach for VLMs. Table 5 presents the calibration analysis for SOTA VLMs, with all methods showing strong calibration in MCQs. For OQs, Qwen2.5-VL significantly outperforms other approaches. The calibration plots for Qwen2.5-VL are shown in Fig. 3. The Section 10.2 of the supplementary material includes deeper calibration analysis.
Table 6: Qwen2.5-VL performance comparison when varying the GT clip length of FALCONBench validation split.
Visual Information MCQs OQs
$\Delta t$ Frames Accuracy Accuracy Score
s # Res. M S T Avg. M S T Avg. M S T Avg.
60 2fps $336\times 616$ 85.0 54.3 73.3 70.9 60.0 57.1 66.6 61.3 3.15 3.23 3.46 3.28
60 64 $476\times 840$ 80.0 62.8 73.3 72.1 70.0 57.1 53.3 60.2 3.60 3.46 2.93 3.38
60 32 $672\times 1204$ 85.0 71.4 86.7 81.0 60.0 60.0 66.7 62.2 3.15 3.54 3.53 3.43
5 2fps $824\times 1462$ 85.0 74.3 80.0 79.8 70.0 62.9 73.3 68.7 3.70 3.71 3.80 3.73
5 5 $824\times 1462$ 85.0 80.0 93.3 86.1 55.0 65.7 66.7 62.5 2.95 3.69 3.53 3.39
Table 7: FALCONEye ablation study in the OQs of FALCONBench validation split. We vary the VLM [Uncaptioned image] (Qwen2.5-VL [4] and LLaVA-Video [52]), the LLM [Uncaptioned image] (GPT4o-mini, GPT4o and, GPTo3-mini), the zoom-in effect ([Uncaptioned image] or fix resolution), the max. number of evaluated candidate clips [Uncaptioned image] and the confidence threshold [Uncaptioned image]. We report the average accuracy (Acc.) and score (Sc.).
FALCONEye LLM VLM Time OQs
N [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] #Inf. #T #Inf.T s Acc. Sc.
[52] 4o-m 45 0.8 7.0 1.5 41 549.1 17.5 1.13
[Uncaptioned image] [4] 4o-m 45 0.8 3.7 1.7K 21.2 229.2 46.6 2.47
[4] 4o 45 0.8 5.1 1.6K 24.4 274.8 48.8 2.56
[4] o3-m 45 0.8 6.5 3.6K 25.6 315.8 50.4 2.63
[4] 4o-m 45 0.8 4.3 1.7K 23.8 258.1 43.0 2.31
[4] 4o-m 25 0.8 3.4 1.7K 19.7 212.7 45.6 2.45
[Uncaptioned image] [4] 4o-m 10 0.8 2.9 1.8K 17.5 188.3 42.6 2.26
[4] 4o-m 45 0.7 4.0 1.7K 18.6 205.4 45.0 2.44
[4] 4o-m 45 0.9 5.8 1.8K 34.9 375.6 49.5 2.64
Zoom-in effect study.
Qwen2.5-VL can process images of any resolution, dynamically converting them into a variable number of visual tokens. This creates a trade-off between number of frames and frame resolution when processing videos within the context window size limit. This trade-off is analyzed in the validation split of FALCONBench by varying the clip length (Tab. 6). The best configuration per clip-level category is selected for FALCONEye.
FALCONEye.
Tab. 7 validates some details of our best configurations FALCONEye [Uncaptioned image] and FALCONEye [Uncaptioned image]. While other configurations are more accurate, we must take into the account efficiency. For LLMs choice, it is important to note that this validation split (70Q) experiment costs: GPT4o-mini \$0.1, GPT4o \$2.0, and GPTo3-mini \$3.5.
7 Conclusions
We present FALCONEye, a novel video agent driven by an efficient exploration algorithm integrating a LLM and a VLM to tackle the challenge of answering open-ended questions in hour-long videos (VAS) with answers located in short temporal windows. We evaluate our meta-architecture using a 7B-size VLM and a cost-efficient LLM (academic resources), and demonstrate that it significantly outperforms sota 7B-size VLMs and similar video agents on FALCON-Bench, a new designed benchmark for the VAS task. Besides, when applied to shorter videos and multiple-choice questions in existing benchmarks, it achieves similar performance on average to top VLMs like GPT-4o.
Supplementary Material
8 FALCON-Bench: more details
8.1 Examples and video source details
The videos of our benchmark were sourced from three different public datasets:
  • SoccerNet [16] – 56 videos were selected from this dataset, each averaging 92.4 minutes in duration. These structured soccer match recordings include 389 questions. The mean GT time interval is 61.9 seconds.
  • MovieChat-1K Films [38] – a total of 140 film clips, each 8 minutes long, were selected from the dataset. Clips from the same film were combined to create 24 film segments, with an average duration of 46.4 minutes each. This subset contains 122 questions, with a mean GT time interval of 22.4 seconds.
  • Walking Tours Dataset [40] – 12 high-resolution videos ($3840\times 2160$) were selected from this dataset, averaging 81.3 minutes. These videos capture city tours from an egocentric perspective and include 65 questions. The mean GT time interval is 31.4 seconds.
Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Figure 4 shows an example question of our benchmark across each dataset.
Figure 4: Falcon-Bench question examples for each dataset.
8.2 Question Categories
To design a benchmark for long VAS tasks, each question should have its answer contained within a single, short clip of the video. Based on this consideration, we defined four question categories:
  • Text Reading (TR): Questions ask about a piece of text that appears at a certain moment in the video.
  • Visual Observation (VO): Questions focus on visual attributes of the items appearing in the video, such as colors, textures, components, or materials.
  • Time Identification (TI): Questions about timestamps on clocks or alarms shown in the video.
  • Object Identification (OI): Questions focus on identifying specific objects within the video.
Figure 5 provides an overview of the number of questions per video type and per category, illustrating the distribution of tasks across the benchmark.
Figure 5: Distribution of questions in Falcon-Bench. The left plot, according to dataset sources: MovieChat-1k, SoccerNet, and WalkingTours. The right plot according to question category: TR, VO, TI, and OI.
8.3 Localization evaluation
FALCONBench requires models to provide an evidence of the answer (short clip) rather than precisely matching the entire clip where the answer is located. To achieve this goal, we leverage The Ground Truth over Union (GToU) metric to compare the predicted and the GT temporal interval. Unlike the commonly used Intersection over Union (IoU) in temporal grounding tasks [13], GToU assigns a score of 1.0 if the predicted interval is entirely within the GT interval, regardless of the degree of overlap. In all other cases, GToU behaves identically to IoU (Figure 6). In mathematical terms,
$\textbf{GToU}=\frac{|GT|}{|GT\cup\text{Pred}|}\cdot\mathbbm{1}_{\{|GT\cap\text{Pred}|>0\}}.$ (2)
Figure 6: Visualization of GToU metric designed to measure the clip localization/retrieval task in which the answer is contained.
8.4 Human experiments
To evaluate human performance on our benchmark, we performed experiments with 10 participants. Each participant answered 10 questions based on 10 different videos (one question per video) and equally spread across the different dataset sources. Each participant answered 5 MCQs and 5 OQs.
Figure 7: Performance metrics across all participants. Figures show accuracy, mean GToU (mGToU), and time spent per question.
Methodology
Participants were seated in front a test computer equipped with two monitors, one for watching the video and another for editing a .json file. They were provided with a structured .json file containing the question details in the following format:
Listing 1: JSON question example provided to volunteers.
"id": "WalkingTours-009",
"question": "Whichcarcompanyhasa
servicecenternexttoToyota?",
"video": "Chiang_Mai.mp4",
"options": null,
"type": "Open-endedquestion",
"answer": null,
"temporal_window": null,
"spent_time": null
Participants were required to fill in the answer and temporal_window fields. About the answer, if the options field is null, the answer was open-ended; otherwise, the answer was a letter in [A, B, C, D]. In the timestamp the participants had to indicate a short video clip where the answer is observed. A supervisor recorded the total time spent for each question, starting when the participant opened the video and stopping when they finished entering both required fields.
Results
The results of the human experiments were analyzed in terms of both individual performance of each participant and the aggregated performance across all 10 participants, in terms of accuracy, mean GToU (mGToU) and spent time to answer. Figure 7 shows individual performance per participant for both MCQs and OQs. Participants 2, 3, 4, and 9 answered correctly all questions. Regarding the spent time, participant 3 was the fastest. Additionally, Table 8 shows the mean accuracy, mGToU, time and score (only for open-ended questions) of the answers across the three video types (MovieChat-1K, SoccerNet, WalkingTours). SoccerNet questions are generally easier to answer correctly, simpler to locate within the video, and quicker to respond to. In contrast, WalkingTours questions are the most challenging overall due to the extended duration and continuous nature of the videos.
Table 8: Mean performance metrics of all participants across the three video types (MovieChat-1K, SoccerNet, WalkingTours).
Dataset Accuracy mGToU Time (s) Score (0-5)
MCQs OQs MCQs OQs MCQs OQs OQs
MovieChat-1k 100.0 84.6 63.6 60.1 118.5 120.2 4.31
SoccerNet 100.0 95.0 79.5 89.2 55.2 67.4 4.80
WalkingTours 76.9 76.5 76.9 58.2 173.1 200.2 4.06
9 FALCONEye: more details
In this section we give further implementation details of our FALCONEye video agent explained in Sec. 3.
9.1 Exploration Algorithm pseudo-code
Algorithm 1 shows the pseudo-code of our FALCONEye exploration algorithm proposed in Sec. 3.2 and which can be visualized in Fig. 2.
Input: Video $\mathcal{V}$, Question $Q$
Output: Answer $\mathcal{A}$, Confidence $p(\mathcal{A})$, $Clip$ $v\in\mathcal{V}$
Hyperparams: $it_{max},dur_{t}$
Stage 1 Pre-processing
$Clips\leftarrow\text{Segment}(\mathcal{V})$
$\mathcal{C}\leftarrow\text{VLM}(Clips)$
$\mathcal{S}\leftarrow\text{LLM}(\mathcal{C})$
$AllAns\leftarrow[\;\;]\;\;;\;\;it\leftarrow 0$
while $it\leq it_{max}$ do
       Stage 2 Reasoning
       $Cand\_Clips\leftarrow\text{LLM}(Q,Captions)$
       $Cand\_Captions\leftarrow Cand\_Clips.Captions$
       while $dur(Cand\_Clips)\geq dur_{t}$ do
             $Ans\leftarrow[\;\;]$
             Stage 3 Evaluation
             foreach $v_{i}^{*}\in Cand\_Clips$ do
                   $\mathcal{A}_{i}^{*},p(\mathcal{A}_{i}^{*})\leftarrow\text{VLM}(Q,v_{i}^{*})$
                   $Ans.append(\{\mathcal{A}_{i}^{*},p(\mathcal{A}_{i}^{*}),v_{i}^{*},c_{i}\})$
                   $it\leftarrow it+1$
             end foreach
            Stage 4 Decision
             $\mathcal{A},Promis\_Clips\leftarrow\text{LLM}(Q,Ans)$
             if $\mathcal{A}\neq None$ then
                   return $\mathcal{A},p(\mathcal{A}),v$
             end if
            $AllAns.append(Ans)$
             $Cand\_Clips\leftarrow\text{Segment}(Promis\_Clips)$
       end while
      $\mathcal{C}.remove(Cand\_Captions)$
end while
$\mathcal{A}\leftarrow\text{LLM}(Q,AllAns)$
return $\mathcal{A},p(\mathcal{A}),v$
Algorithm 1 FALCONEye Exploration Algorithm
9.2 Prompts
Figure 8 shows the prompts sent to the LLM during the different stages of our FALCONEye exploration algorithm. Regarding the stages explained in Sec. 3.2: Summary generation from stage 1 Pre-processing, Select candidate clips from captions from stage 2 Reasoning, Final answer or keep exploring and Select candidate clips as promising clips to keep exploring from stage 4 Decision, and Return final answer when the maximum number of evaluated candidate clips is reached.
Figure 8: LLM prompts used through FALCONEye algorithm.
10 Additional Experiments
10.1 FALCON-Bench
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we further evaluated VLMs under a simplified VAS setup. Specifically, we extracted 1-minute clips centered within the ground truth intervals to ensure that the answers were contained within these shorter segments (Table 9). Qwen2.5-VL achieved the best accuracy and score in OQs, and LLaVA-Video in MCQs. However, even after reducing the search space from an hour-long video to just one minute, the performance remains relatively low. To address this, we took an additional step and tested the VLMs by providing a single frame that contains the answer, which is always within the ground truth interval defined by FALCONBench (Table 10). These results represent the maximum performance our FALCONEye agent could achieve with each VLM, assuming the ground truth frame is the optimal frame for answering the question. In this evaluation, GPT-4o achieved the highest performance, followed by Qwen2.5-VL.
Table 9: Model performance on FALCONBench test split with GT 1min-length clips containing the answer. We report average accuracy for MCQs and OQs across MovieChat (M), SoccerNet (S), WalkingTours (T), and overall (Avg.).
MCQs OQs
Acc. Accuracy Score (0-5)
Model Name Avg. M S T Avg. M S T Avg.
Baselines
Full Mark 100 100 100 100 100 5.00 5.00 5.00 5.00
Random 25.0 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Open-Source VLMs with Multi-Image Support
LLaVa-v1.5 33.3 6.86 4.23 4.00 5.03 0.49 0.37 0.40 0.42
LLaVa-v1.6 41.9 9.80 23.4 8.00 13.7 0.58 1.33 0.58 0.83
mPLUG-Owl3 [47] 57.6 39.2 21.1 26.0 28.7 2.24 1.20 1.58 1.67
LLaVA-OV [17] 72.2 45.0 28.5 46.0 39.8 2.46 1.57 2.42 2.15
Qwen2.5-VL 75.9 64.7 52.2 64.0 60.3 3.49 3.08 3.26 3.27
Open-Source VLMs designed for Videos
VideoChat2-HD [21] 27.2 18.6 12.1 22.0 17.5 1.24 0.67 1.34 1.08
Video-LLAVA [22] 34.0 19.6 6.49 18.0 14.6 1.20 0.46 1.14 0.93
LLaVA-Video [19] 79.2 61.7 44.0 52.0 52.6 3.31 2.38 2.74 2.81
Open-Source VLMs specific for long videos
MovieChat-OV [38] 44.5 5.88 18.3 14.0 12.7 0.49 1.02 0.82 0.77
Apollo [54] 71.3 50.9 31.3 48.0 43.4 2.77 1.79 2.62 2.39
Table 10: Model performance on the test set providing the model with GT frames containing the answer. We report average accuracy for MCQs and OQs across MovieChat (M), SoccerNet (S), WalkingTours (T), and overall (Avg.).
MCQs OQs
Acc. Accuracy Score
Model Name Avg. M S T Avg. M S T Avg.
Baselines
Full Mark 100 100 100 100 100 5.00 5.00 5.00 5.00
Random 25.0 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Proprietary Long-Context LLMs
GPT-4o-mini (LR) 75.3 61.7 28.5 42.0 44.0 3.18 1.52 2.28 2.32
GPT-4o-mini (HR) 88.4 70.5 57.9 70.0 66.1 3.58 3.21 3.78 3.52
GPT-4o (LR) 83.7 78.0 47.7 62.0 52.5 3.94 2.63 3.32 3.29
GPT-4o (HR) 94.1 80.3 66.9 72.0 73.0 4.06 3.67 3.82 3.85
Open-Source VLMs for images
LLaVa-v1.5 56.7 40.1 24.5 34.0 32.8 2.34 1.38 1.96 1.89
LlaVa-v1.6 78.5 53.9 48.5 62.0 54.8 3.08 2.51 3.42 3.24
mPLUG-Owl3 [47] 82.9 59.8 53.1 58.0 65.2 3.25 2.89 3.18 3.10
LLaVA-OV [17] 84.4 40.1 10.4 54.0 34.8 2.02 0.53 2.80 1.78
Qwen2.5-VL [] 84.1 71.5 62.9 78.0 70.8 3.80 3.33 4.02 3.72
LLaMa3.2M [1] 85.9 69.6 64.9 70.0 68.1 3.68 3.27 3.70 3.55
10.2 VLMs Calibration
To measure VLM calibration, we adopt the Reliability Diagrams [12]. These diagrams group all the predictions in bins according to their confidence and measure the gap between confidence and accuracy per each bin. Specifically, we split the $N$ predictions in $M=10$ bins according to their confidence and, for each bin $B_{m}$, we compute its count $N_{m}$ average confidence $C_{m}$ and its average accuracy $A_{m}$. From these diagrams, we may compute the Average Calibration Error (ACE) as,
$\text{ACE}=\frac{1}{M^{+}}\sum_{m=1}^{M}|C_{m}-A_{m}|,\;\;\text{MCE}=\max_{m}|C_{m}-A_{m}|$ (3)
where $M^{+}$ is the number of non-empty bins [30]. We compute Calibration Count (CC), which quantifies the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE). For example, the Calibration Count at threshold 0.9 is computed as,
${\text{CC}@0.9=\frac{N_{10}}{N}\left(1-|C_{10}-A_{10}|\right)}.$ (4)
Confidence metrics-.
As detailed in Sec.3.2, estimating answer confidence in OQs requires a metric to average token probabilities along the response (see Figure 9).
We evaluated various confidence aggregation methods, specifically likelihood, average, and geometric average, as done in similar calibration studies with LLMs [26]. Looking first at the reliability diagrams (Figure 10 (a-f)), we discard likelihood as a suitable confidence metric. The distribution of confidence values is highly spread out, with many answers assigned extremely low confidence. This occurs because likelihood multiplies the probability of all tokens, thus longer answers tend to receive lower confidence scores, making it unreliable for calibration.
Comparing the average and geometric average, the reliability diagrams do not show major differences between them. However, the calibration metrics in Table 11 indicate that geometric average results in lower Brier Score (BS), MCE, and ACE for both LLaVA-Video and Qwen2.5-VL, suggesting slightly better calibration performance.
Figure 9: Given a question and a set of frames, FALCONEye leverages the answer outputted by the VLM but, and its confidence (geometric average through all output tokens probabilities).
Table 11: Calibration metrics comparison with GT 1min-length clips in the open-ended (OQs) questions test split of FALCONBench. Lower values for BS, MCE and ACE, are better.
Model Name $\Delta t$ F Likelihood Average Geom. Average
s # BS MCE ACE BS MCE ACE BS MCE ACE
LLaVA-Video 60 32 0.23 0.34 0.17 0.24 0.37 0.22 0.22 0.31 0.16
LLaVA-OV 60 32 0.26 0.92 0.41 0.28 0.39 0.29 0.26 0.59 0.29
Qwen2.5-VL 60 2fps 0.19 0.31 0.15 0.26 0.45 0.25 0.24 0.38 0.27
Refer to caption Refer to caption Refer to caption Refer to caption
(a) LLaVA-Video, Likelihood. (b) Qwen2.5-VL-2fps, Likelihood. (g) LLaVA-Video, MCQs (h) LLaVA-Video, OQs
Refer to caption Refer to caption Refer to caption Refer to caption
(c) LLaVA-Video, Average. (d) Qwen2.5-VL-2fps, Average. (i) LLaVA-OV, MCQs (j) LLaVA-OV, OQs
Refer to caption Refer to caption Refer to caption Refer to caption
(e) LLaVA-Video, Geom. avg (f) Qwen2.5-VL-2fps, Geom. avg. (k) Qwen2.5-VL-2fps, MCQs (l) Qwen2.5-VL-2fps, OQs
Figure 10: (a-f) Calibration plots for different probabilities aggregation metrics in the open-question evaluation scenario with the GT 1min clips of the FALCONBench test split. We focus on both LlaVa-Video and Qwen2.5-VL models. (g-l) calibration plots when testing the VLMs with the GT 1min-length clips of the FALCONBench test split for both MCQs (left) and OQs (right) with the geometric average aggregation metric.
Figure 11: Answer comparison between GPT4o and FALCONEye for three example questions, showing the frame that contain the answer.
Models-.
Regarding model selection, we compare reliability diagrams for LLaVA-Video, LLaVA-OneVision, and Qwen2.5-VL, evaluated on both MCQs and OQs. For MCQs, all models show strong calibration, with low calibration error, specially for high confidence values (Fig. 10 (g-l)). However, for OQs, the distribution of high-confidence answers differs significantly. Both LLaVA-Video and LLaVA-OneVision have a much smaller proportion of answers with 0.9 confidence, whereas Qwen2.5-VL not only produces a higher number of high-confidence answers but also demonstrates extremely low calibration error for those predictions.
10.3 FALCONEye vs GPT4o
Figure 11 show three example questions and compare the responses from GPT4o and FALCONEye.